Naresh walked 2 km 35 m in the morning and 1 km 7 m in the evening. How much distance did he walk in all?

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

Distance walked by Naresh in the morning = 2 km 35 m
= [(2 + 35 /1000)] km
= 2.035 km
Distance walked by him in the evening = 1 km 7 m
= [(1 + 7 / 1000)] km
= 1.007 km
Total distance walked by Naresh is
=2.035 + 1.007
= 3.042
∴Total distance walked by Naresh is 3.042 km

Thanks
Shraddha Singh
3 years ago
Dear Student,Distance walked by Naresh in the morning = 2 km 35 m                                                                            = [(2 + 35 /1000)] km= 2.035 kmDistance walked by him in the evening = 1 km 7 m                                                                     = [(1 + 7 / 1000)] km= 1.007 km★Total distance walked by Naresh is=2.035 + 1.007                                                                  = 3.042